Using an appropriate measure,compare the variation of the following data set
Weights of newborn (kg):3.2,3.3,3.4,3.5,2.7,2.8,3.1
Heights of female(inches):70,72,70,67,65,80,71
Answer #1

Answer:

Using an appropriate measure,compare the variation of the following data set
Weights of newborn (kg):3.2,3.3,3.4,3.5,2.7,2.8,3.1
Heights of female(inches):70,72,70,67,65,80,71

\bar x =\frac{\sum x}{n}

s =\sqrt{ \frac{\sum x^2- (\sum x)^2 /n)}{n-1}}

coefficient of variation (CV) = (sd/mean)*100

Descriptive statistics

Weights

Heights

n

7

7

mean

3.1429

70.7143

sample standard deviation

0.2992

4.7509

minimum

2.7

65

maximum

3.5

80

coefficient of variation (CV)

9.52%

6.72%

coefficient of variation for weights is 9.52% and heights is 6.72%.

Variation in weights is larger than variation in heights.